# Causal Inference in Longitudinal Data under Unknown Interference

### v4

@@ A.2 Proof for Proposition 1
${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]={\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}{\rm E}\,\left[\frac{\mathbf{1}\{A_{j}^{(T)}=a^{(T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{P\left(A_{j}^{(T)}=a^{(T)}\mid\mathbf{V}_{j}^{(T)}\right)}\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}{\rm E}\,\left[Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(T)},A_{j}^{(T)}=a^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}{\rm E}\,\left[Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\right)}\right]=\dots$
@@
The $\dots$ part iterates the same step from period $T-1$ to period $1$. The first equality uses the definition of the potential outcome, while the fourth equality uses the law of total expectation. The fifth equality holds because of sequential exchaneagibility. Therefore, for any $d\in\mathcal{D}$,
$\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]=\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu\left(\big{\{}Y_{i}\big{\}}_{i\in\mathcal{N}};d\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]$
$=$ $\frac{1}{N}\sum_{j=1}^{N}\frac{1}{|\Omega_{j}(d)|}\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\right)}\right]$
@@
In a saturated MSM, each of the parameters is a linear combination of the AMRs thus can also be identified from data.
A.3 Proof for Theorem 1
We start from the case where $e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)$ is known. Consider the first term in the expression of $\hat{\beta}(d)$, namely, $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$. We can establish the following lemma:
Lemma A.1
@@ A.3 Proof for Theorem 1
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]={\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]{\rm E}\,\left[\frac{f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]$
@@
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]={\rm E}\,\left[{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\mid\mathbf{V}_{i}^{(T)}\right]\right]$
@@
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right]\right]$
@@
$=$ $\sum_{\mathbf{a}_{i}^{(1:T)}\in\mathcal{A}^{(1:T)}}\sum_{\mathbf{a}_{j}^{(1:T)}\in\mathcal{A}^{(1:T)}}f\left(\mathbf{a}_{i}^{(1:T)}\right)f\left(\mathbf{a}_{j}^{(1:T)}\right)={\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right]{\rm E}\,\left[\frac{f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}\right].$
In our framework, $\mathbf{L}_{i}^{(T)}$ and $\mathbf{L}_{j}^{(T)}$ are dependent because both are functions of $\mathbf{A}_{\mathcal{N}}^{(1:T)}$. As a result, for any function $f(\cdot)$, $f\left(\mathbf{A}_{i}^{(1:T)}\right)$ and $f\left(\mathbf{A}_{j}^{(1:T)}\right)$ are generally dependent as well. Nevertheless, Lemma A.1 suggests that once normalized by the weights, $\frac{1}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}$ and $\frac{1}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)}\right)}$, they become uncorrelated. Due to positivity and the binary nature of $A_{i}^{(T)}$, $w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$ is uniformly bounded. Then, using Lemma A.1, we can see that
@@
We further define $\mathbf{U}_{\mathcal{H}}(d)=(\dots,U_{i}(d),\dots)^{\prime}_{i\in\mathcal{H}}$ and $\mathbf{U}_{\mathcal{H}}^{(p,\dagger)}(d)$ similarly. For any subset ~$\tilde{\mathcal{H}}\subset\mathcal{N}_{N}$ such that $\min_{i\in\mathcal{H},j\in\tilde{\mathcal{H}}}d_{ij}\geq p$ and $\mathbf{U}_{\tilde{\mathcal{H}}}^{(p,\ddagger)}(d)$ defined on it, it is straightforward to see that ${\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]=0$ if $p>2*d$. Then, for any $f_{h}\in\mathcal{L}_{h}$ and $f_{\tilde{h}}\in\mathcal{L}_{\tilde{h}}$, we know that $\Bigg{|}{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]\Bigg{|}\leq 2||f_{h}||_{\infty}||f_{\tilde{h}}||_{\infty}$ for $p\leq 2*d$, and for $p>2*d$,
@@
$\leq$ $2\Big{|}||f_{\tilde{h}}||_{\infty}{\rm E}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right)-f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right)\right]\Big{|}+2\Big{|}||f_{h}||_{\infty}{\rm E}\,\left[f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)-f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]\Big{|}$
$\leq$ $2||f_{\tilde{h}}||_{\infty}hL_{h}\theta_{N,\lfloor\frac{p}{2}\rfloor}(d)+2||f_{h}||_{\infty}\tilde{h}L_{\tilde{h}}\theta_{N,\lfloor\frac{p}{2}\rfloor}(d).$
@@
A.4 Proof for Theorem 2
First, note that $\frac{\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}}{N}\to\Sigma_{A}$ as $N\to\infty$ by Lemma A.1 and standard theory. Define $\mathbf{Q}^{\prime}_{i}(d)=\sum_{k=1}^{N}\frac{\mathbf{1}\{i\in\Omega_{k}(d)\}w_{k}\mathbf{m}\left(\mathbf{A}_{k}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{k}^{(1:T)}\right)}{|\Omega_{k}(d)|}$. Then, $\hat{\boldsymbol{\varepsilon}}_{Yi}(d)=\mathbf{Z}_{i}(d)Y_{i}-\mathbf{Q}^{\prime}_{i}(d)\hat{\boldsymbol{\beta}}(d)$. It is straightforward to verify that $\frac{1}{N}\sum_{i=1}^{N}\mathbf{Q}^{\prime}_{i}(d)=\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$, which implies that $\boldsymbol{\beta}(d)=\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Q}^{\prime}_{i}(d)\right]\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]\right)$ and $\hat{\boldsymbol{\beta}}(d)=\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Q}^{\prime}_{i}(d)\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\right)$. Next, we prove the following lemma concerning the terms $\{\mathbf{Z}_{i}(d)Y_{i}\}_{i\in\mathcal{N}_{N}}$ and $\{\mathbf{Q}_{i}(d)\}_{i\in\mathcal{N}_{N}}$:
Lemma A.2

### v5

@@ B.1 Proof for Proposition 1
${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]={\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))};\mathbf{V}_{j}^{(1:(T-1))}\right)}{\rm E}\,\left[\frac{\mathbf{1}\{A_{j}^{(T)}=a^{(T)}\}Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)}{P\left(A_{j}^{(T)}=a^{(T)}\mid\mathbf{V}_{j}^{(T)}\right)}\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))};\mathbf{V}_{j}^{(1:(T-1))}\right)}{\rm E}\,\left[Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(T)},A_{j}^{(T)}=a^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))};\mathbf{V}_{j}^{(1:(T-1))}\right)}{\rm E}\,\left[Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\mid\mathbf{V}_{j}^{(T)}\right]\right]$
$=$ ${\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))}\}Y_{i}\left(a^{(T)},\mathbf{A}_{j}^{(1:(T-1))},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:(T-1))}=\mathbf{a}^{(1:(T-1))};\mathbf{V}_{j}^{(1:(T-1))}\right)}\right]=\dots$
@@
The $\dots$ part iterates the same step from period T1T-1 to period $1$. The first equality uses the definition of the potential outcome, while the fourth equality uses the law of total expectation. The fifth equality holds because of sequential exchaneagibility. Therefore, for any $d\in\mathcal{D}$,
$\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]=\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu\left(\big{\{}Y_{i}\big{\}}_{i\in\mathcal{N}};d\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]$
$=$ $\frac{1}{N}\sum_{j=1}^{N}\frac{1}{|\Omega_{j}(d)|}\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}Y_{i}}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]$
@@
In a saturated MSM, each of the parameters is a linear combination of the AMRs thus can also be identified from data. We can similarly prove that
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\left[{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]\right]=\mu^{*}\left(\mathbf{a}^{(1:T)};d\right).$
B.2 Proof for Theorem 1
We start from the case where $e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ is known. Consider the first term in the expression of $\hat{\beta}(d)$, namely, $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$. We can establish the following lemma:
Lemma B.1
@@ B.2 Proof for Theorem 1
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]{\rm E}\,\left[\frac{f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]$
@@
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]={\rm E}\,\left[{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\mid\mathbf{V}_{i}^{(T)}\right]\right]$
@@
${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]$
$=$ ${\rm E}\,\left[{\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\mid\mathbf{V}_{i}^{(T)},\mathbf{V}_{j}^{(T)}\right]\right]$
@@
$=$ $\sum_{\mathbf{a}_{i}^{(1:T)}\in\mathcal{A}^{(1:T)}}\sum_{\mathbf{a}_{j}^{(1:T)}\in\mathcal{A}^{(1:T)}}f\left(\mathbf{a}_{i}^{(1:T)}\right)f\left(\mathbf{a}_{j}^{(1:T)}\right)$
$=$ ${\rm E}\,\left[\frac{f\left(\mathbf{A}_{i}^{(1:T)}\right)}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right]{\rm E}\,\left[\frac{f\left(\mathbf{A}_{j}^{(1:T)}\right)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right].$
In our framework, $\mathbf{L}_{i}^{(T)}$ and $\mathbf{L}_{j}^{(T)}$ are dependent because both are functions of $\mathbf{A}_{\mathcal{N}}^{(1:T)}$. As a result, for any function $f(\cdot)$, $f\left(\mathbf{A}_{i}^{(1:T)}\right)$ and $f\left(\mathbf{A}_{j}^{(1:T)}\right)$ are generally dependent as well. Nevertheless, Lemma B.1 suggests that once normalized by the weights, $\frac{1}{e\left(\mathbf{A}_{i}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}$ and $\frac{1}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{A}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}$, they become uncorrelated. Due to positivity and the binary nature of $A_{i}^{(T)}$, $w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$ is uniformly bounded. Then, using Lemma B.1, we can see that
@@
[…]
@@
[…]
@@
To establish the convergence of $\hat{\boldsymbol{\beta}}(d)$ to $\boldsymbol{\beta}^{*}(d)$ defined in Section A.1, we impose additional restrictions to bound the dependence in $Y_{i}$ induced by the dependence in $\mathbf{F}_{i}^{(1:T)}$. For any two $\sigma$-fields $\sigma$ and $\tilde{\sigma}$, define
$\alpha\left(\sigma,\tilde{\sigma}\right)=\sup_{\mathcal{A}\in\sigma,\mathcal{B}\in\tilde{\sigma}}\bigg{|}\mathbbm{Cov}\left(\mathbf{1}\{\mathcal{A}\},\mathbf{1}\{\mathcal{B}\}\right)\bigg{|},$
where $\mathbbm{Cov}(\cdot)$ is defined using $\mathbb{E}[\cdot]$. For any two sets $\mathcal{H},\tilde{\mathcal{H}}\in\mathcal{N}_{N}$ such that $d\left(\mathcal{H},\tilde{\mathcal{H}}\right)>p$, we further define the strong mixing coefficients as
[…]
We then impose the following assumption:
Assumption B.4 (Strong mixing of stochastic factors)
$\sup_{N}\alpha_{N,p}\rightarrow 0,$
as $p\rightarrow\infty$.
Assumption B.4 indicates that any dependence in the stochastic factors declines to zero as the proximity between two units increases.
For any set $\mathcal{H}$, ${\rm E}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right)\mid\mathbf{F}_{\mathcal{H}}^{(1:T)}\right]$ is a random variable defined on $\sigma\left(\Big{\{}\mathbf{F}_{i}^{(1:T)}\Big{\}}_{i\in\mathcal{H}}\right)$, which is bounded by $\bigg{|}\bigg{|}{\rm E}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right)\mid\mathbf{F}_{\mathcal{H}}^{(1:T)}\right]\bigg{|}\bigg{|}_{\infty}$. This quantity is finite since $Y_{i}$ is uniformly bounded. Then, under Assumption B.4, when p>2dp>2*d, we have
$\mathbbm{Cov}\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]=\mathbb{E}\left[{\rm Cov}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\right]\right]$
$+\mathbbm{Cov}\left[{\rm E}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right)\mid\mathbf{F}_{\mathcal{H}}^{(1:T)}\right],{\rm E}\,\left[f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\mid\mathbf{F}_{\tilde{\mathcal{H}}}^{(1:T)}\right]\right]$
$\leq 4\bigg{|}\bigg{|}{\rm E}\,\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}^{(\lfloor\frac{p}{2}\rfloor,\dagger)}(d)\right)\mid\mathbf{F}_{\mathcal{H}}^{(1:T)}\right]\bigg{|}\bigg{|}_{\infty}\bigg{|}\bigg{|}{\rm E}\,\left[f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}^{(\lfloor\frac{p}{2}\rfloor,\ddagger)}(d)\right)\mid\mathbf{F}_{\tilde{\mathcal{H}}}^{(1:T)}\right]\bigg{|}\bigg{|}_{\infty}\alpha_{N,p}=\tilde{\alpha}_{N,p}\to 0.$
The last inequality uses Theorem 9 in Prakasa Rao (2009). Consequently, we can similarly show that
$\Bigg{|}\mathbbm{Cov}\left[f_{h}\left(\mathbf{U}_{\mathcal{H}}(d)\right),f_{\tilde{h}}\left(\mathbf{U}_{\tilde{\mathcal{H}}}(d)\right)\right]\Bigg{|}$
[…]
Assumption 3.3 in Kojevnikov, Marmer and Song (2021) is thus satisfied, and consistency and asymptotic normality follow.
B.3 Proof for Theorem 2
First, note that $\frac{\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}}{N}\to\Sigma_{A}$ as $N\to\infty$ by Lemma B.1 and standard theory. Define $\mathbf{Q}^{\prime}_{i}(d)=\sum_{k=1}^{N}\frac{\mathbf{1}\{i\in\Omega_{k}(d)\}w_{k}\mathbf{m}\left(\mathbf{A}_{k}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{k}^{(1:T)}\right)}{|\Omega_{k}(d)|}$. Then, $\hat{\boldsymbol{\varepsilon}}_{Yi}(d)=\mathbf{Z}_{i}(d)Y_{i}-\mathbf{Q}^{\prime}_{i}(d)\hat{\boldsymbol{\beta}}(d)$. It is straightforward to verify that $\frac{1}{N}\sum_{i=1}^{N}\mathbf{Q}^{\prime}_{i}(d)=\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$, which implies that $\boldsymbol{\beta}(d)=\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Q}^{\prime}_{i}(d)\right]\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}{\rm E}\,\left[\mathbf{Z}_{i}(d)Y_{i}\right]\right)$ and $\hat{\boldsymbol{\beta}}(d)=\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Q}^{\prime}_{i}(d)\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\right)$. Next, we prove the following lemma concerning the terms $\{\mathbf{Z}_{i}(d)Y_{i}\}_{i\in\mathcal{N}_{N}}$ and $\{\mathbf{Q}_{i}(d)\}_{i\in\mathcal{N}_{N}}$:
Lemma B.2